Problem: The total luminous flux falling onto the floor and the walls from the electrical

bulb that is fixed to the ceiling of the room is 314 lm. Find the luminous intensity of the

bulb if the luminous flux is the same in all directions.


1
Expert's answer
2022-01-10T09:08:00-0500

"I=\frac \Phi \Omega=\frac{314}{4\cdot 3.14}=25~kd."